# MultiBanana: A Challenging Benchmark for Multi-Reference Text-to-Image Generation

### v1

@@ A.2 Details on Image Filtering
A.3 Details on Category Classification
To construct multi-reference image editing tasks, each reference image had to be accurately categorized. For example, in portrait transformation tasks that modify a person’s hairstyle or makeup, both reference images must contain a person. Similarly, in replacement tasks involving humans or objects, the corresponding entities must be present in the reference images, and for background replacement tasks, background scenes must be determined in the reference images. To satisfy these task-dependent requirements, we performed hierarchical image classification on a large combined set of real and synthetic images.
For synthetic data, attributes can be specified directly during generation, eliminating the need for additional classification. In contrast, real images exhibit substantial variability, requiring precise categorization to determine whether they can serve as valid reference images. To address this, we designed a structured annotation protocol using Gemini that consists of three classification processes for the human, object, and background components of an image.
Human Image Classification  Human classification begins with identifying how many people appear in the image. If there is precisely one person, the model additionally classifies attributes such as gender, makeup, hairstyle distinctiveness, pose clarity, and facial expression characteristics. The output follows a fixed template format, allowing consistent extraction of appearance features relevant to multi-reference generation.
@@ A.3 Details on Category Classification
Background Image Classification  Background classification decomposes the scene into realism, tonal uniformity, and lighting strength. Realism distinguishes photographic scenes from stylized or synthetic ones; tone indicates whether the overall color distribution is uniform or varied; and lighting strength identifies whether a clear directional key light is present. Ambiguous lighting is conservatively labeled as absent to maintain annotation consistency.
A.4 Details on Task Construction
@@
The editing instructions were generated using Gemini. First, we manually selected suitable reference categories for each task, for instance, the person category for the hairstyle modification task. Next, we randomly sampled images from each category and presented them to Gemini to produce corresponding editing instructions. For example, in editing tasks that require object transformations, we first prompted the model to determine which reference image’s object should be modified and which reference image’s attributes (e.g., color) should be used as the target. Afterward, several instruction examples were provided to Gemini, and we asked Gemini to generate new instructions following the same syntax. After that, editing instructions that would cause visual breakdowns were removed by scoring them with Gemini. Additionally, the difficulty of editing instructions was also assessed by Gemini, and those judged easy, such as cases without domain differences, were removed. Finally, the editing instructions were manually verified not to cause image breakdowns or result in overly simple edits, and assigned the remaining samples to the appropriate difficulty categories (cross-domain, scale and viewpoint differences, rare concept, and multilingual text rendering).
B Further Statistics for MultiBanana
@@
The left of Figure 9 shows the number of tasks by each reference count. The two-reference tasks contain more samples than the other tasks because they include multiple task types. The right of Figure 9 shows the breakdown of the two-reference tasks. It contains eleven tasks considered in the prior work [xia2025dreamomni2]: subject addition, subject replacement, background change, color modification, material modification, pose modification, hairstyle modification, makeup modification, tone transformation, style transfer, and text correction. Each task is ensured to include at least 6% of the editing samples, corresponding to roughly 60 samples, which is larger than the number of samples in xia2025dreamomni2.
Figure 10: Qualitative example of “Hard” category of ImgEdit. These tasks are close to being fully solvable by advanced models such as Nano Banana.
Figure 11: Qualitative example of DreamOmni2 benchmark. These tasks are close to being fully solvable by advanced models such as Nano Banana.
C Further Comparison with Prior Benchmarks
Table 1 shows that existing benchmarks do not provide systematic evaluation across diverse multi-reference conditions, support only a limited number of references, and fail to adequately account for important factors such as differences and compatibility among reference images. To illustrate that existing benchmarks are almost solved by state-of-the-art models such as Nano Banana [google2025nanobanana] and GPT-Image-1 [openai2025gpt4oimage], we evaluate the image generation capabilities of these models using ImgEdit [ye2025imgedit], the latest benchmark for image editing, and DreamOmni2 [xia2025dreamomni2], the state-of-the-art benchmark for multi-reference image generation. The quantitative results show that recent closed-source, state-of-the-art models achieve consistently high scores, indicating that these benchmarks are nearing saturation and may soon be unable to meaningfully distinguish among high-performing models (Table 4 and Table 5). The evaluation results are taken from the official GitHub repository of ImgEdit [^1] https://github.com/PKU-YuanGroup/ImgEdit?tab=readme-ov-file and the results reported in the DreamOmni2 paper. Our qualitative results suggest that, even in the additional “Hard” category of ImgEdit, most tasks are already close to being fully solvable by advanced models such as Nano Banana (Figure 10). A similar tendency is observed for DreamOmni2 benchmark as well (Figure 11).
Taken together, these findings indicate that current benchmarks for image editing and multi-reference image generation are approaching their limits in evaluating cutting-edge models. There is a clear need for more challenging benchmarks that can more effectively capture the capabilities and failure modes of the latest generation of image models.
@@ C Further Comparison with Prior Benchmarks
Table 4: Comparison results of different models on ImgEdit-Bench [ye2025imgedit]. “Overall” is computed by averaging the scores across all task types. GPT-4.1 [openai2023gpt4] is used for evaluation. The results show that recent closed-source, state-of-the-art models achieve substantially higher scores, suggesting that the benchmark may be approaching its ceiling in distinguishing high-end models. Evaluation results excluding Nano Banana are taken from the official GitHub repository of ye2025imgedit.
@@
MagicBrush [zhang2023magicbrush] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.83
AnyEdit [yu2025anyedit] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.45
OmniGen2 [wu2025omnigen2] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 3.44
Kontext-dev [labs2025flux1kontext] 3.83 3.65 2.27 4.45 3.17 3.98 4.55 3.35 3.71
Nano Banana [google2025nanobanana] 3.63 4.66 3.73 4.69 4.71 4.60 4.47 4.10 4.37
GPT-Image-1 [openai2025gpt4oimage] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.20
Table 5: Quantitative comparison of multimodal instruction-based editing and generation in DreamOmni2 Benchmark [xia2025dreamomni2]. Both tasks are evaluated using Gemini [geminiteam2023gemini] and Doubao [bytedance2025doubao]. This result shows that recent closed-source, state-of-the-art models achieve substantially higher scores, indicating that the benchmark may be reaching its ceiling for distinguishing among high-end models. We refer to evaluation results from xia2025dreamomni2.
@@
Omnigen2 [wu2025omnigen2] 0.2195 0.2927 0.0427 0.0793 0.2083 0.2500 0.1000 0.0778
Kontext [labs2025flux1kontext] 0.0488 0.1220 0.0183 0.0122 0.2500 0.3750 0.0556 0.1222
Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport] 0.2683 0.2927 0.0488 0.1159 0.1250 0.2917 0.1111 0.1556
DreamOmni2 [xia2025dreamomni2] 0.5854 0.6585 0.5854 0.6280 0.5833 0.6667 0.5778 0.6333
Nano Banana [google2025nanobanana] 0.6829 0.7073 0.6463 0.5488 0.5000 0.5417 0.5556 0.5488
GPT-Image-1 [openai2025gpt4oimage] 0.6829 0.7805 0.7195 0.7439 0.6250 0.6250 0.6889 0.6333
D Further Discussions for AI and Human Evaluation
The correlation between AI and human evaluation has been shown in prior works [na2024boost, wu2024boosting, furuta2024improving, oshima2025inference]. Consequently, using VLMs is becoming a standard method for the automatic evaluation of generated images [ye2025imgedit, wu2025omnigen2, xia2025dreamomni2]. We used VLMs to evaluate the images along five metrics: Instruction Alignment, Reference Consistency, Background Subject Match, Physical Realism, and Visual Quality. The overall score is calculated as a weighted average of these metrics. The first two metrics represent the fundamental components of reference-based generation, while the last three metrics break down the overall quality. Therefore, the overall score is calculated as $\mathrm{avg}(\mathrm{Instruction\,Alignment}+\mathrm{Reference\,Consistency}+\mathrm{avg}(\mathrm{Background\,Subject\,Match}+\mathrm{Physical\,Realism}+\mathrm{Visual\,Quality}))$.
To demonstrate the consistency of our evaluation metrics, we examined the correlation between different VLMs. We use the evaluated results for Nano Banana and GPT-Image-1, and examine the correlation coefficient between GPT’s and Gemini’s evaluation score. To ensure the reliability of the correlation coefficient, we divided the evaluation results into ten subsets and computed the average correlation across them. Table 6 shows that the evaluated scores exhibit positive correlations, which demonstrates that our evaluation criteria are well-defined and consistent.
Table 6: Correlation coefficients between GPT’s and Gemini’s evaluated scores. $\pm$ shows 95% confidence interval.
@@ D Further Discussions for AI and Human Evaluation
E Detailed Results
E.1 Further Discussion of Per-Task Evaluation
Figure 12, Figure 13, and Figure 14 show the scores of each model across five evaluation metrics for each task, evaluated by GPT, Gemini, and their average, respectively. Taking the weighted average across five metrics, Figure 15 shows the overall score of each model for single and two-reference tasks, evaluated by GPT, Gemini, and their average, respectively. For multi-reference tasks, Table 8, Table 9, and Table 10 show the overall score, evaluated by GPT, Gemini, and their average, respectively. Overall, we observed a substantial gap in Instruction Alignment and Reference Consistency between the closed-source models and the other open-source models. In particular, GPT-Image-1 achieves significantly superior performance in Instruction Alignment, especially in style and text modification tasks in the two-reference task. Meanwhile, in Reference Consistency, Nano Banana achieves the highest scores in the single-reference task. Nano Banana also performs on par with GPT-Image-1 in Reference Consistency, across background modification, subject replacement, subject addition, and makeup tasks, which indicates its strong ability to leverage reference images, particularly when the number of references is small. Additionally, Qwen-Image-Edit-2509, the most advanced open-source model, performs worse than the other models in background modification tasks in Background Subject Match and Visual Quality. This result suggests that it struggles with background modification.
Figure 14: Average scores of each model across evaluation metrics for each task by GPT and Gemini. The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
Figure 15: Overall scores of each model for single and two-reference tasks by GPT, Gemini, and their average.
Table 8: Detailed per-task overall scores for the multi-reference tasks, evaluated by GPT. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@ E.1 Further Discussion of Per-Task Evaluation
Table 9: Detailed per-task overall scores for the multi-reference tasks, evaluated by Gemini. The local, global, back, and object columns under X correspond to X-1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
Table 10: Detailed per-task overall scores for each multi-reference task, evaluated by the average of GPT and Gemini. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
Figure 16: Overall scores for each difficult reference combination across models. Darker colors represent the average score of tasks that include the corresponding combination, while lighter colors indicate the average score of tasks that do not include it.
Figure 17: Qualitative example for 4-Object Tasks.
Figure 18: Qualitative example for 3-Object + Background Tasks.
Figure 19: Qualitative example for 3-Object + Local Tasks.
Figure 20: Qualitative example for 3-Object + Global Tasks.
Figure 21: Qualitative example for Tasks with 8 references.
E.2 Further Discussion of the Effect of the Number of References
According to the per-metric scores in Figure 14 and the overall scores in Table 10, all models exhibit decreasing performance as the number of reference images increases. For Instruction Alignment and Reference Consistency, GPT-Image-1 and Nano Banana achieve relatively high scores, whereas the open-source model, DreamOmni2, tends to approach the minimum score of 1 as the number of references increases. In contrast, for quality-related metrics, Background Subject Match, Physical Realism, and Visual Quality, the DreamOmni2 maintains high scores even when more reference images are provided. As discussed in Section 4, this is because closed-source models prioritize adherence to the references and editing instructions over visual quality, while open-source models show a tendency to preserve visual quality but often ignore the references.
In the finer-grained task-level evaluation, we found that even as the number of reference images increased, the X–1 Object + Global task did not exhibit score degradation in Background Subject Match, Physical Realism, or Visual Quality. This is because the task requires modifying the image to a specified style; therefore, even if the reference images come from different domains, which may degrade the visual quality, the final output converges to a unified style.
In the results of Qwen-Image-Edit-2509, the model is trained to handle up to four reference images; therefore, when provided with five or more references, the input becomes out-of-distribution. This causes the model to produce perceptually invalid outputs, and then the scores approach the minimum value of 1.
E.3 Further Discussion of Difficult Reference Combination

### v2

@@ B.2 Image Filtering
Figure 10: Example of “Hard” category of ImgEdit. These tasks are almost fully solvable by advanced models such as Nano Banana.
Figure 11: Example of DreamOmni2 benchmark. These tasks are almost fully solvable by advanced models such as Nano Banana.
B.3 Category Classification
To construct multi-reference image editing tasks, each reference image had to be accurately categorized. For example, in portrait transformation tasks that modify a person’s hairstyle or makeup, both reference images must contain a person. Similarly, in replacement tasks involving humans or objects, the corresponding entities must be present in the reference images, and in background replacement tasks, the background scenes must be identified in the reference images. To satisfy these task-dependent requirements, we performed hierarchical image classification on a large combined set of real and synthetic images.
For synthetic data, attributes can be specified directly in the prompt during generation, eliminating the need for additional classification. In contrast, real images exhibit substantial variability, requiring precise categorization to determine whether they can serve as valid reference images. To address this, we designed a structured annotation protocol using Gemini that consists of three classification processes for the human, object, and background components of an image.
Human Image Classification  Human classification begins with identifying how many people appear in the image. If there is precisely one person, the model additionally classifies attributes such as gender, makeup, hairstyle distinctiveness, pose clarity, and facial expression characteristics. The output follows a fixed template, enabling consistent extraction of appearance features relevant to multi-reference generation.
@@ B.3 Category Classification
Background Image Classification  Background classification is conducted from three viewpoints: realism, tonal uniformity, and lighting strength. Realism distinguishes photographic scenes from stylized or synthetic ones; tone indicates whether the overall color distribution is uniform or varied; and lighting strength identifies whether a clear directional key light is present.
B.4 Task Construction
@@
The editing instructions were generated using Gemini. First, we manually selected suitable reference categories for each task, for instance, the person category for the hairstyle modification task. Next, we randomly sampled images from each category and presented them to Gemini to produce corresponding editing instructions. For example, in editing tasks that require object transformations, we first prompted the model to determine which reference image’s object should be modified and which reference image’s attributes (e.g., color) should be used as the target. Afterward, several instruction examples were provided to Gemini, and we asked Gemini to generate new instructions following the same syntax. Afterward, editing instructions that could cause visual breakdowns were removed after being scored with Gemini. Additionally, Gemini assessed the difficulty of editing instructions, and those judged easy, such as cases without domain differences, were removed. Finally, the editing instructions were manually verified not to cause image breakdowns or result in overly simple edits, and assigned the remaining samples to the appropriate difficulty categories (cross-domain, scale and viewpoint differences, rare concept, and multilingual text rendering).
C Further Statistics for MultiBanana
@@
The left of Figure 11 shows the number of tasks by each reference count. The two-reference tasks contain more samples than the other tasks because they include multiple task types. The right of Figure 11 shows the breakdown of the two-reference tasks. It contains eleven tasks considered in the prior work [83]: subject addition, subject replacement, background change, color modification, material modification, pose modification, hairstyle modification, makeup modification, tone transformation, style transfer, and text correction. Each task is guaranteed to include at least 6% of the editing samples, corresponding to roughly 60 samples, which exceeds the number of samples in Xia et al. [83].
@@
Table 7: Comparison results of different models on ImgEdit-Bench [91]. “Overall” is computed by averaging the scores across all task types. GPT-4.1 [54] is used for evaluation. The results show that recent closed-source, state-of-the-art models achieve substantially higher scores, suggesting that the benchmark may be approaching its ceiling in distinguishing high-end models. Evaluation results excluding Nano Banana are taken from the official GitHub repository of Ye et al. [91].
@@
MagicBrush [93] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.83
AnyEdit [92] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.45
OmniGen2 [80] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 3.44
Kontext-dev [41] 3.83 3.65 2.27 4.45 3.17 3.98 4.55 3.35 3.71
Nano Banana [22] 3.63 4.66 3.73 4.69 4.71 4.60 4.47 4.10 4.37
GPT-Image-1 [56] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.20
Table 8: Quantitative comparison of multimodal instruction-based editing and generation in DreamOmni2 Benchmark [83]. Both tasks are evaluated using Gemini [19] and Doubao [5]. This result shows that recent closed-source, state-of-the-art models achieve substantially higher scores, suggesting that the benchmark may be approaching its ceiling in distinguishing among high-end models. We refer to evaluation results from Xia et al. [83].
@@
Omnigen2 [80] 0.2195 0.2927 0.0427 0.0793 0.2083 0.2500 0.1000 0.0778
Kontext [41] 0.0488 0.1220 0.0183 0.0122 0.2500 0.3750 0.0556 0.1222
Qwen-Image-Edit-2509 [79] 0.2683 0.2927 0.0488 0.1159 0.1250 0.2917 0.1111 0.1556
DreamOmni2 [83] 0.5854 0.6585 0.5854 0.6280 0.5833 0.6667 0.5778 0.6333
Nano Banana [22] 0.6829 0.7073 0.6463 0.5488 0.5000 0.5417 0.5556 0.5488
GPT-Image-1 [56] 0.6829 0.7805 0.7195 0.7439 0.6250 0.6250 0.6889 0.6333
D Comparison with Prior Benchmarks
Table 1 shows that existing benchmarks do not provide systematic evaluation across diverse multi-reference conditions, support only a limited number of references, and fail to adequately account for heterogeneity among reference images. To illustrate that existing benchmarks are almost solved by state-of-the-art models such as Nano Banana [22] and GPT-Image-1 [56], we evaluate the image generation capabilities of these models using ImgEdit [91], the latest benchmark for image editing, and DreamOmni2 [83], the state-of-the-art benchmark for multi-reference image generation. The quantitative results show that recent closed-source, state-of-the-art models achieve consistently high scores, indicating that these benchmarks are nearing saturation and may soon be unable to meaningfully distinguish among high-performing models (Table 8 and Table 8). The evaluation results are taken from the official GitHub repository of ImgEdit [^1] https://github.com/PKU-YuanGroup/ImgEdit?tab=readme-ov-file and the results reported in the DreamOmni2 paper. Qualitative results suggest that, even in the additional “Hard” category of ImgEdit, most tasks are already solvable by advanced models such as Nano Banana (Figure 11). A similar tendency is observed for DreamOmni2 benchmark as well (Figure 11).
Taken together, these findings indicate that current benchmarks for image editing and multi-reference image generation are approaching their limits in evaluating cutting-edge models. There is a clear need for more challenging benchmarks that better capture the capabilities and failure modes of the latest generation of image models.
E Reliability and Cost of VLM Judges
As described in Section 3.4, we employ VLM judges as a time- and cost-efficient proxy for human evaluation. In Section 4.4, we demonstrate that the VLM judge correlates strongly with human ratings, supporting our usage of VLM. Here, we provide a more detailed analysis of their reliability and the associated cost considerations.
E.1 Cross-VLM Correlation
To demonstrate the consistency of our evaluation metrics, we examined correlations among different VLMs. We use the evaluated results for Nano Banana and GPT-Image-1 and examine the correlation between GPT’s and Gemini’s evaluation scores. To assess the reliability of the correlation coefficient, we partitioned the evaluation results into 10 subsets and computed the mean correlation across them. Table 9 shows that the evaluated scores exhibit positive correlations, which demonstrates that our evaluation criteria are well-defined and consistent.
Table 9: Correlation between GPT’s and Gemini’s evaluated scores. $\pm$ shows 95% confidence interval.
@@ E.1 Cross-VLM Correlation
E.2 Self-Consistency Analysis
We measure the standard error of the VLM judge scores across three different random seeds in Table 10. On the 1–10 rating scale, the observed standard errors are below 0.1, and we do not observe an increasing trend in standard error as the number of references increases, indicating that the VLM judge provides self-consistent evaluations for multi-reference image generation.
Table 10: Standard error of total scores evaluated by the average of GPT and Gemini, for GPT-Image-1 and Nano Banana.
Model 3-ref 4-ref 5-ref 6-ref 7-ref 8-ref
GPT-Image-1 .0234.0234 .0138.0138 .0139.0139 .0153.0153 .0343.0343 .0244.0244
Nano Banana .0260.0260 .0423.0423 .0033.0033 .0089.0089 .0062.0062 .0254.0254
E.3 Sensitivity Analysis
@@
E.4 Cost of VLM Judge
Regarding evaluation costs, a single model evaluation costs approximately \$44 with GPT-5 and \$29 with Gemini 2.5 in January 2026. As an API-free alternative, we confirmed that Qwen3-VL [2] can be reliable enough (see Table 4).
F Detailed Results
F.1 Per-Task Evaluation
Figure 15, Figure 16, and Figure 17 show the scores of each model across five evaluation metrics for each task, evaluated by GPT, Gemini, and their average, respectively. Taking the weighted average across five metrics, Figure 18 shows the total score of each model for single and two-reference tasks, evaluated by GPT, Gemini, and their average, respectively. For multi-reference tasks, Table 12, Table 13, and Table 14 show the total score, evaluated by GPT, Gemini, and their average, respectively.
Overall, we observed a substantial gap in Instruction Alignment and Reference Consistency between the closed-source models and the other open-source models. In particular, GPT-Image-1 achieves significantly superior performance in Instruction Alignment, especially in style and text modification tasks in the two-reference task. Meanwhile, in Reference Consistency, Nano Banana achieves the highest scores in the single-reference task. Nano Banana also performs on par with GPT-Image-1 in Reference Consistency across background modification, subject replacement, subject addition, and makeup tasks, indicating its strong ability to leverage reference images, particularly when the number of references is small. On the other hand, Nano Banana performs worse than the other models in background modification tasks in Background Subject Match, Physical Realism, and Visual Quality. This result suggests that Nano Banana struggles with background modification.
@@ F.1 Per-Task Evaluation
Table 12: Detailed per-task total scores for the multi-reference tasks, evaluated by GPT. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
Table 13: Detailed per-task total scores for the multi-reference tasks, evaluated by Gemini. The local, global, back, and object columns under X correspond to X-1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
Table 14: Detailed per-task total scores for each multi-reference task, evaluated by the average of GPT and Gemini. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
F.2 The Effect of the Number of References
According to the per-metric scores in Figure 17 and the total scores in Table 14, all models exhibit decreasing performance as the number of reference images increases. For Instruction Alignment and Reference Consistency, GPT-Image-1 and Nano Banana achieve relatively high scores, whereas the open-source model, DreamOmni2, tends to approach the minimum score of 1 as the number of references increases. In contrast, for quality-related metrics, Background Subject Match, Physical Realism, and Visual Quality, the DreamOmni2 maintains high scores even when more reference images are provided. As discussed in Section 4, this is because closed-source models prioritize adherence to the references and editing instructions over visual quality, while open-source models show a tendency to preserve visual quality but often ignore the references.
In the finer-grained task-level evaluation, we found that, even as the number of reference images increased, the X–1 Object + Global task did not show severe decreases in Background Subject Match, Physical Realism, or Visual Quality scores. The task requires modifying the image to a specified style; therefore, even if the reference images come from different domains, which may degrade the visual quality, the final output often converges to a unified style.
In the results of Qwen-Image-Edit-2509, the model may be trained to handle up to three reference images; therefore, when provided with four or more references, the input becomes out-of-distribution. This causes the model to produce perceptually invalid outputs, and then the scores approach the minimum value of 1 (see Figure 29).
F.3 Difficult Reference Combination